### v2

# SAM Decoding: Speculative Decoding via Suffix Automaton

Abstract
Large Language Models (LLMs) have revolutionized natural language processing by unifying tasks into text generation, yet their large parameter sizes and autoregressive nature limit inference speed. SAM-Decoding addresses this by introducing a novel retrieval-based speculative decoding method that uses a suffix automaton for efficient and accurate draft generation. Unlike n-gram matching used by the existing method, SAM-Decoding finds the exact longest suffix match in generating text and text corpuss, achieving an average time complexity of $O(1)$ per generation step. Meanwhile, based on the exact match length, it is designed as an approach that can be combined with existing methods, allowing SAM-Decoding to adaptively select a draft generation strategy based on the matching length, thus increasing the inference speed of the LLM. When combined with Token Recycling, evaluations show SAM-Decoding outperforms existing model-free methods, achieving a speedup of 2.27×2.27$\times$2.27 × over autoregressive decoding on Spec-Bench. When combined with EAGLE2, it reaches a speedup of 2.49×2.49$\times$2.49 ×, surpassing all current approaches. Our code is available at ours repository.
SAM Decoding: Speculative Decoding via Suffix Automaton
Yuxuan Hu1,2, Ke Wang1,2, Xiaokang Zhang1,2, Fanjin Zhang4 Cuiping Li1,3Hong Chen1,3Jing Zhang1,3thanks: Corresponding author. 1School of Information, Renmin University of China, Beijing, China 2Key Laboratory of Data Engineering and Knowledge Engineering, Beijing, China 3Engineering Research Center of Database and Business Intelligence, Beijing, China 4Knowledge Engineering Group, Tsinghua University, Beijing, China
1 Introduction
The Transformer-based Large Language Model Brown et al. (2020); Dubey et al. (2024); Yang et al. (2024) has become a foundation model in the field of natural language processing. Through the large language model (LLM), all natural language processing tasks are unified as text generation tasks, which greatly improves the convenience of solving natural language processing tasks. While LLM has shown good performance in various tasks and is easy to use, the large-scale parameters and text generation-based paradigm contained in LLM have brought challenges to the practical deployment, especially in terms of the inference speed of the LLM. LLM generates tokens step by step in an autoregressive manner. At each round of generation, LLM generates one new token based on the tokens that has been generated. In addition, for the GPUs used for deploy LLM, this process requires reading the huge number of parameters contained in the LLM from the High-Bandwise Memory (HBM) to the on-chip memory, but these parameters only operate on a small number of parameters corresponding to the new token, resulting in the inability to fully utilize the parallel capabilities of the modern GPU, thereby limiting the inference speed of the LLM.
To address the above problem, one effective strategy is to enhance the arithmetic intensity of the large language model (LLM) generation process, which involves reducing the number of generation rounds while increasing the number of tokens generated in each round. Building upon this concept, the method of speculative decoding has been introduced and has spurred a series of follow-up studies Leviathan et al. (2023); Miao et al. (2024); Cai et al. (2024); Fu et al. (2024). Rather than generating a single token per step, these approaches employ a draft model to predict token sequence as draft. The predicted sequence are then verified against the LLM, and only the accurate tokens are adopted.
Most speculative decoding techniques employ various neural networks, such as smaller language models (LLMs) or embedding layers, as draft models, which we categorize as model-based methods. Although model-based speculative methods perform well on all kinds of tasks, when we consider some special tasks such as summarization, rewriting, there is a simpler type of speculative decoding methods that exhibit superior performance, which we call retrieval-based speculative decoding methods. These methods involve retrieving potential subsequent tokens from previously generated content or a text database during each generation round. Unlike model-based approaches, retrieval-based methods do not necessitate the training of draft models or the use of GPUs for inference.
Figure 1: Overview of SAM-Decoding’s workflow. SAM-Decoding includes two automaton: static SAM and dynamic SAM. Among them, the static SAM is built based on a given text corpus, while the dynamic SAM is built based on generating text and gradually expanded during the generation process. When drafting, the suffix automaton matches the suffixes of the generating text and retrieves the draft from the text corpus and the generating text respectively according to the matching position. Meanwhile, we also combined an auxiliary speculative decoding algorithm (Aux Method) to deal with scenarios where retrieval is not applicable. we select the draft to be used from the three candidate drafts based on the match length. After that, the draft will be verified by LLM to obtain accepted tokens. Using these accepted tokens, we finally extend the dynamic SAM and generating text for the next round of generation.
After examining the current retrieval-based methods, we identify several limitations. Firstly, these approaches typically rely on a single retrieval source. For instance, PLD Saxena (2023) focuses solely on prompts, whereas REST He et al. (2024) is limited to text corpus. Secondly, there are limitations in the retrieval techniques they adopt; PLD relies on n-gram matching, which is fast for context matching but cannot be extended to a wider range of textual repositories due to its poor theoretical complexity, while REST is based on suffix arrays, which has better complexity but cannot be applied to continually generating text. Lastly, not every step in the generation is suitable for the retrieval method; in a few generation steps, a large number of tokens can be accepted based on PLD and REST, but in the rest of the generation steps there is no acceleration.
To address the limitations of the retrieval-based approach mentioned earlier, this paper introduces SAM-Decoding, an innovative retrieval-based speculative decoding technique. Unlike existing approaches that rely on n-gram matching, SAM-Decoding employs a suffix automaton (SAM) to solve the exact longest suffix match problem for the generating text within both the historical generated text and the text corpus. This process yields more accurate match positions and exact match lengths compared to n-gram matching. Moreover, during the text generation phase, the suffix automaton effectively captures the relationships between adjacent suffixes, ensuring that the average time complexity for identifying the longest suffix match at each step remains at $O(1)$. This represents a significant improvement over the theoretical computational efficiency of n-gram matching. Meanwhile, the length of the longest suffix match serves as an indicator of the alignment between the generating text and retrieval sources. By exploiting this metric, our approach can integrate existing speculative decoding methods as an auxiliary and dynamically decide whether to generate a draft based on the matched content, or utilize an auxiliary decoding method. This adaptability leads to a notable increase in the overall speed of the text generation process.
Specifically, SAM-Decoding first constructs a static suffix automaton for the existing text corpus. For each request, SAM-Decoding then builds a dynamic suffix automaton based on the input prompt. The suffix automaton contains a series of nodes, each of which corresponds to several substrings in the context (text corpus, prompt and generating text). We record the earliest corresponding position of each node in the text. This information facilitates the generation of drafts by directly mapping nodes in the automaton to indices in the context. We also maintain the corresponding nodes of the generating text in both the static and dynamic automaton. Therefore, in each round of generation, we can directly retrieve the draft from the context according to the position recorded in the node, and filter the draft according to the matching length. After generation, as new tokens are generated, for the static automaton, we transition nodes of SAM based on new tokens, and for the dynamic automaton, we first expand the structure to accommodate the new tokens before transitioning nodes of SAM accordingly.
Extensive evaluations across a range of tasks indicate that the SAM-Decoding can achieve results that are competitive with existing state-of-the-art (SOTA) approaches in all tasks. Particularly in tasks where retrieval-based methods are applicable, SAM-Decoding outperforms all existing methods. On Spec-Bench, the latest benchmark for speculative decoding, SAM-Decoding combined with Token Recycling Luo et al. (2024) achieves a speedup of 2.26 times compared to autoregressive decoding, outperforming all model-free baselines. Additionally, when SAM-Decoding is combined with EAGLE2 Li et al. (2024a), the SOTA model-based speculative decoding method, it achieves a speedup of 2.49 times compared to autoregressive decoding, outperforming all existing model-based and model-free methods.
2 Background
2.1 Suffix Automaton
Suffix Automaton is an efficient data structure for representing the substring index of a given string, which allows fast substring retrieval. The time complexity of constructing a suffix automaton is $O(L)$, where $L$ is the length of the string and it can be constructed incrementally.
A suffix automaton contains a series of nodes and two types of state transfer edges, extension edges and suffix link edges. A node in the automaton represents a state and corresponds to all substrings that have the same ending position in the string. Meanwhile, extension edges are standard edges that represent a possible extension of the current substring by appending a new character, while suffix link edges create a path that allows the automaton to quickly jump to states representing shorter suffixes of the current substring.
Based on the two types of transfer edges, for a progressively generated token sequence, we can find the longest suffix that matches the sequence in a suffix automaton at each step of the generation with an average $O(1)$ time complexity.
2.2 Speculative Decoding
Given the model input $x=(x_{1},x_{2},\ldots,x_{t})$, an LLM generates a new token $x_{t+1}$ at each generation step autoregressively. The key idea of speculative decoding is to utilize a lightweight draft model to generate multiple candidate tokens quickly, i.e., $x_{\text{draft}}=(x_{t+1},x_{t+2},\ldots,x_{t+n})$, and then the target LLM simultaneously evaluates these candidates and accept those aligned with the output distribution of the LLM, i.e., $x_{\text{accept}}=(x_{t+1},x_{t+2},\ldots,x_{t+m})$.
In the above, we assume that the draft is a sequence of tokens. Recent works proposed to generate a candidate token tree via a tree mask in the attention module to make the target LLM simultaneously evaluate multiple branches of this token tree, thereby increasing the acceptance length of the draft model.
3 SAM-Decoding
In this section, we introduce our proposed method, SAM-Decoding. SAM-Decoding is a retrieval-based speculative decoding method designed to address three key limitations in existing retrieval-based speculative methods: (1) The use of insufficient retrieval sources. (2) The employment of inefficient retrieval methods and restrictions on $n$-gram matching lengths. (3) Subpar performance outside specialized domains (e.g., summarization and RAG tasks).
To tackle the first two limitations, SAM-Decoding leverages suffix automaton, which significantly enhances the coverage of retrieved corpus and the efficiency of the retrieval process while allowing for more flexible matching lengths. In what follows, we detail how SAM-Decoding can be integrated with both model-free and model-based methods. By utilizing the precise matching information provided by the suffix automaton, our method not only overcomes the third limitation but also ensures consistent performance improvements across a wide range of tasks.
3.1 Suffix Automaton Construction
SAM-Decoding facilitates rapid draft generation by leveraging suffix automaton. Unlike traditional methods that rely on a single retrieval source, our approach utilizes both a text corpus and the generating text itself as retrieval sources. To accommodate these dual sources, we construct two distinct types of suffix automaton: a static suffix automaton and a dynamic suffix automaton. For the text corpus, we pre-build a static suffix automaton offline. During the inference phase, state matching is performed exclusively on this static automaton. Conversely, for the generating text, we create a dynamic suffix automaton that incrementally expands as the text generation progresses. State matching is also carry out concurrently with this expansion.
A suffix automaton can be constructed in linear time using Blumer’s algorithm Blumer et al. (1984). Since the suffix automaton is designed for a single reference string, when dealing with multiple strings in the text corpus, we concatenate them into a single long string using special symbols, such as an End-of-Sentence (EOS) token. We then construct a static suffix automaton for this concatenated string.
At the same time, we also made several modifications to the suffix automaton to better suit draft generation. Specifically, at each generation step, the generating text corresponds to a node in the automaton, representing the longest suffix match. We need to identify the matching position in either the text corpus or the generating text. Based on this position, we can use the subsequent tokens as potential drafts. To achieve this, at every node of the suffix automaton, we record the earliest position of all substrings corresponding to that node in the reference string. We define this position as min_endpos. By using the min_endpos, we can directly map node to the position in the reference string. The construction process of the suffix automaton is detailed in Appendix A.1.
3.2 Drafting with Suffix Automaton
After constructing the suffix automaton, it can be utilized as a draft model to generate draft at each step of the generation process. Specifically, let the current input to the LLM be denoted as $x=(x_{1},x_{2},\ldots,x_{t})$, with the predicted next token being $x_{t+1}$. We define the suffix automaton as $S$ and its associated reference text as $T$. The state within the suffix automaton corresponding to the sequence $x$ is denoted as $s_{t}$. The transition to the next state is performed based on the newly generated token $x_{t+1}$ and the current state $s_{t}$:
$s_{t+1}$ $=\text{Transfer}(S,x_{t+1},s_{t}).$
Subsequently, we extract $n$ consecutive tokens from the reference text $T$ to form a draft, using the min_endpos value stored in node corresponding to state $s_{t+1}$. If the min_endpos value in $s_{t+1}$ is $p_{t+1}$, then the draft $d$ is defined as:
$\text{\rm d}=T[p_{t+1}:p_{t+1}+n],$
where $d$ represents the generated draft and $n$ denotes the length of the draft.
In practical use, we utilize two types of automaton: a static automaton for the text corpus and a dynamic automaton for generating text. To distinguish between the two, we explicitly record the attribute $l_{t+1}$ for the state $s_{t+1}$. This attribute represents the longest suffix length matched by the generating text in the automaton. This helps us determine whether the draft should be generated using the static suffix automaton or the dynamic suffix automaton. Our experimental findings indicate that drafts generated from the dynamic automaton often outperform those from the static text corpus. Consequently, we prioritize drafts from the dynamic automaton. Specifically, let $l_{1}$ and $l_{2}$ represent the matching lengths of the static automaton and the dynamic automaton, respectively. We use the draft provided by the static automaton only if $l_{1}>l_{2}+l_{\text{bias}}$; otherwise, we use the draft provided by the dynamic automaton. Here, $l_{\text{bias}}$ is a predefined constant greater than 0.
We illustrate the state transfer process of the suffix automaton in Algorithm 1. Using amortized analysis, we can prove that the average complexity of state transfer is $O(1)$, while the worst-case time complexity is $O(L)$, where $L$ represents the length of the generating text. The detailed proof is provided in Appendix A.2. Reviewing existing methods, PLD uses $n$-grams and performs a brute-force search for matches in the prompt, resulting in a time complexity of $O(n^{2}L)$. REST also employs $n$-grams but searches for matches in the text using suffix arrays, leading to a time complexity of $O(n^{2}\log L)$. Here, $n$ represents the predefined maximum matching length, and $L$ represents the length of the prompt and the total text in the corpus, respectively. In comparison, our proposed model (SAM-Decoding) has a lower time complexity. Additionally, SAM-Decoding can find the exact longest suffix match without any limit on the matching length. This means that SAM-Decoding can generate drafts both faster and more accurately.
Algorithm 1 State Transfer of Suffix Automaton
  function Transfer
     Input: suffix automaton $S$, next token $t$, current state $s$, current matching length $l$
     while $s\neq S.\text{root}$ and $t\notin s.\text{next}$ do
        $s=s.\text{link}$
        $l=s.\text{length}$
     end while
     if $t\in s.\text{next}$ then
        $s=s.\text{next}[t]$
        $l=l+1$
     else
        $l=0$
     end if
     return $s$, $l$
  end function
3.3 Update of Suffix Automaton
After the draft is generated, we verify the draft using the large language model (LLM) and accept the correct tokens from the draft. Let the accepted tokens be denoted as $x_{\text{accept}}=(x_{t+1},x_{t+2},\ldots,x_{t+m})$. We then update the state of the suffix automaton based on these accepted tokens. For the static suffix automaton, we simply transfer the matching states according to Algorithm 1:
$s_{t+i}=\text{Transfer}(S,s_{t+i-1},x_{t+i}),\;i\in\{1,2,...,m\}.$
For the dynamic suffix automaton, we first expand its state based on the accepted tokens and then transfer the matching state. Let $S_{t}$ denote the dynamic suffix automaton corresponding to the generated text $(x_{1},x_{2},\ldots,x_{t})$. The process is as follows:
$S_{t+i}$ $=\text{Expand}(S_{t+i-1},x_{t+i}),\;i\in\{1,2,...,m\},$
$s_{t+i}$ $=\text{Transfer}(S_{t+m},s_{t+i-1},x_{t+i}),\;i\in\{1,2,...,m\},$
where the process of expanding the suffix automaton is detailed in Appendix A.1.
3.4 Adaptive Draft Selection
A straightforward idea is that the length of the suffix match can indicate the quality of the draft produced by the automaton. A longer match suggests that more tokens from the draft are likely to be acceptable, whereas a shorter match implies that the draft contains less useful information. Consider that besides the retrieval-based speculative decoding approach, there exist various alternative speculative decoding techniques. Consequently, if the retrieval-based method fails to produce a satisfactory draft, we can resort to these alternatives to enhance the draft’s effectiveness.
To implement this, we concurrently employ an additional speculative decoding technique alongside the suffix automaton. During each iteration of the generation process, we adaptively select between the draft offered by the automaton and that provided by the supplementary speculative decoding method, based on the match length of the generated text within the automaton. The adaptive draft selection procedure is detailed in Algorithm 2, where $l_{\rm threshold}$ denotes a predetermined constant that determines whether to utilize the automaton or the auxiliary speculative decoding for draft generation. In our study, we incorporate two cutting-edge speculative decoding methods as auxiliary options: the model-free Token Recycling and the model-based EAGLE. Among them, Token Recycling maintains an adjacency list of the top $k$ probable next tokens for each token. During each generation step, a draft tree is built using this adjacency list via breadth-first search, with the list being continuously updated according to the latest generated tokens. Conversely, EAGLE constructs a draft tree in each generation cycle by leveraging a Transformer decoder layer that integrates the hidden states output by the LLM.
Algorithm 2 Adaptive Draft Selection
  function Drafting
     Input: static suffix automaton $S_{\rm static}$, dynamic suffix automaton $S_{\rm dyn}$, auxiliary speculative method $A$, next token $t$
     $d_{1},l_{1}=\text{Drafting}(S_{\rm static},t)$
     $d_{2},l_{2}=\text{Drafting}(S_{\rm dyn},t)$
     if $l_{1}>l_{2}+l_{\rm bias}$ then
        $d=d_{1}$
        $l=l_{1}$
     else
        $d=d_{2}$
        $l=l_{2}$
     end if
     if $l\geq l_{\rm threshold}$ then
        return $d$
     else
        return $\text{Drafting}(A,t)$
     end if
  end function
4 Experiments
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
Vicuna-7B Lookahead 1.29×1.29$\times$1.29 × 1.03×1.03$\times$1.03 × 1.22×1.$2$ 1.07×1.07$\times$1.07 × 1.41×1.41$\times$1.41 × 1.13×1.13$\times$1.13 × $1.62$ 44.37 1.19×1.19$\times$1.19 ×
REST 1.59×1.59$\times$1.59 × 1.22×1.$2$ 1.27×1.27$\times$1.27 × 1.45×1.45$\times$1.45 × 1.38×1.38$\times$1.38 × 1.36×1.36$\times$1.36 × $1.63$ 51.34 1.38×1.38$\times$1.38 ×
PLD 1.60×1.60$\times$1.60 × 0.95×0.95$\times$0.95 × 2.44×2.$4$ 1.18×1.18$\times$1.18 × 1.59×1.59$\times$1.59 × 1.72×1.72$\times$1.72 × $1.75$ 59.02 1.56×1.56$\times$1.56 ×
SAM-Decoding 1.75$\times$ 0.99$\times$ 2.54$\times$ 1.24$\times$ 1.74$\times$ 1.71$\times$ 1.89 62.26 1.65$\times$
Token Recycling 1.92$\times$ 1.61$\times$ 1.96$\times$ 1.71$\times$ 2.16$\times$ 1.68$\times$ 2.83 69.65 1.84$\times$
SAM-Decoding[T] 2.48$\times$ 1.73$\times$ 2.86$\times$ 1.98$\times$ 2.44$\times$ 2.14$\times$ 3.03 85.73 2.27$\times$
EAGLE 2.63$\times$ 1.92$\times$ 2.28$\times$ 2.11$\times$ 2.64$\times$ 1.95$\times$ 3.57 85.71 2.27$\times$
SAM-Decoding[E] 2.78$\times$ 1.88$\times$ 2.65$\times$ 2.12$\times$ 2.57$\times$ 2.10$\times$ 3.77 88.99 2.35$\times$
EAGLE2 2.87$\times$ 1.92$\times$ 2.33$\times$ 2.20$\times$ 2.88$\times$ 2.03$\times$ 4.36 90.14 2.38$\times$
SAM-Decoding[E2] 3.02$\times$ 1.89$\times$ 2.76$\times$ 2.19$\times$ 2.83$\times$ 2.23$\times$ 4.61 94.17 2.49$\times$
Table 1: Speedup of SAM-Decoding compared to the baselines on Spec-Bench.
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Vicuna-7B PLD 1.65 59.04 1.52×1.52$\times$1.52 × 1.42 50.62 1.27×1.27$\times$1.27 × 2.03 44.11 1.29×1.29$\times$1.29 ×
Token Recycling 2.78 75.44 1.94×1.94$\times$1.94 × 2.83 79.22 1.99×1.$9$ 2.88 66.17 1.93×1.93$\times$1.93 ×
SAM-Decoding[T] 2.94 95.08 2.45×2.45$\times$2.45 × 2.87 94.50 2.37×2.37$\times$2.37 × 3.23 87.93 2.57×2.57$\times$2.57 ×
EAGLE 4.10 103.39 2.66×2.$6$ 4.17 119.31 3.00×3.$0$ 3.44 72.14 2.11×2.$1$
SAM-Decoding[E] 4.03 110.16 2.83×2.83$\times$2.83 × 3.85 114.92 2.87×2.87$\times$2.87 × 4.05 93.17 2.72×2.72$\times$2.72 ×
EAGLE2 5.12 128.19 3.30×3.30$\times$3.30 × 5.29 135.24 3.40×3.40$\times$3.40 × 4.15 82.61 2.41×2.41$\times$2.41 ×
SAM-Decoding[E2] 4.79 127.03 3.27×3.27$\times$3.27 × 4.50 117.06 2.94×2.94$\times$2.94 × 4.75 96.60 2.81×2.81$\times$2.81 ×
Table 2: Speedup of SAM-Decoding compared to the baselines on HumanEval, MBPP, and HAGRID.
In this section, we first introduce our experimental setup, then present the experimental results, and finally present the ablation experiments.
Models and tasks. We conducted experiments on Vicuna-7B-v1.3 Zheng et al. (2023). We evaluated SAM-Decoding on Spec-Bench Xia et al. (2024), HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and HARGID Kamalloo et al. (2023). Spec-Bench is a comprehensive benchmark designed for assessing Speculative Decoding methods across diverse scenarios. It is based on six commonly used datasets, MT-Bench Zheng et al. (2023), WMT14 DE-EN, CNN/Daily Mail Nallapati et al. (2016), Natural Question Kwiatkowski et al. (2019), GSM8K Cobbe et al. (2021), and DPR Karpukhin et al. (2020), including six aspects: Multi-turn Conversation, Translation, Summarization, Question Answering, Mathmatical Reasoning, and Retrieval-augmented Generation. In addition, MBPP, HumanEval, and HARGID are used to evaluate the speed of decoding methods in code generation tasks and contextual question answering tasks, respectively.
Baselines. We considered the following baseline methods, including the model-based method EAGLE Li et al. (2024b, a), the model-free methods Token Recycling Luo et al. (2024) and Lookahead Decoding Fu et al. (2024), and the retrieval-based methods PLD Saxena (2023) and REST He et al. (2024).
Metrics. We evaluated speculative decoding methods from the following aspects Li et al. (2024b)
  • Speedup Ratio: The wall-time speedup ratio of speculative decoding methods compared to autoregressive generation methods.
  • Mean Accepted Tokens: The average number of tokens accepted per generation step.
  • Throughput: The average number of tokens generated per second.
Experiment setup. We conducted experiments on a server equipped with a 20-core CPU and a single NVIDIA RTX A6000 GPU (48GB). The experiments were implemented using PyTorch 2.3.0 and CUDA 12.1. For the models, we used the float16 data type and applied greedy decoding with a batch size of 1. Regarding hyperparameters, unless otherwise specified, $l_{\rm bias}$ and $l_{\rm threshold}$ were set to 5, and the size of the draft generated by the automaton was set to 40 tokens. For the auxiliary speculative decoding algorithms, we used the default configurations as described in their respective original papers. One exception was made for Token Recycling: instead of the draft tree structure used in the original paper, we employed a draft tree with a depth of 6 and containing 60 nodes. In addition, for the vicuna model, we use the vicuna template to convert prompt and output into the chat form, and for the Llama3 model, we use the template provided in the EAGLE2 repository.
For SAM-Decoding, we constructed a static suffix automaton based on the Stanford-alpaca, python-code-instruction-18k, and GSK8k. To enhance our model, we incorporated two auxiliary approaches: the model-free method Token Recycling and the model-based method EAGLE. Here, SAM-Decoding[T], SAM-Decoding[E], and SAM-Decoding[E2] denote the combinations of our base model with Token Recycling, EAGLE, and EAGLE2, respectively.
Experiment Results. Experiment results on Spec-Bench when using Vicuna-7B-v1.3 are shown in Table 1. In the Spec-Bench, which consists of six tasks, multiturn conversation, summarization, and retrieval-augmented generation were identified as particularly amenable to retrieval techniques. The results indicate that integrating SAM-Decoding into these tasks led to notable speed improvements. Specifically, for Token Recycling, the speedup ratio for the three tasks rose from 1.92×1.92$\times$1.92 ×, 1.96×1.96$\times$1.96 ×, and 1.68×1.68$\times$1.68 × to 2.48×2.48$\times$2.48 ×, 2.86×2.86$\times$2.86 ×, and 2.14×2.14$\times$2.14 ×, respectively. With EAGLE, the speedup ratio increased from 2.63×2.63$\times$2.63 ×, 2.28×2.28$\times$2.28 ×, and 1.95×1.95$\times$1.95 × to 2.78×2.78$\times$2.78 ×, 2.65×2.65$\times$2.65 ×, and 2.10×2.10$\times$2.10 ×, respectively. For EAGLE2, the speedup ratios went up from 2.87×2.87$\times$2.87 ×, 2.33×2.$3$, and 2.03×2.03$\times$2.03 × to 3.02×3.02$\times$3.02 ×, 2.76×2.76$\times$2.76 ×, and 2.23×2.23$\times$2.23 ×, respectively. However, for the remaining tasks, the drafts produced by the retrieval methods did not significantly aid speculative decoding and instead introduced extra computational costs, leading to slight decreases in performance. Despite this, when SAM-Decoding was combined with these existing methods, there was an overall increase in speedup ratios: from 2.83×2.83$\times$2.83 × to 2.98×2.98$\times$2.98 × for Token Recycling, from 3.57×3.57$\times$3.57 × to 3.77×3.$7$ for EAGLE, and from 2.38×2.38$\times$2.38 × to 2.49×2.49$\times$2.49 × for EAGLE2. It is noteworthy that in the summarization task, the model-free approach (SAM-Decoding[T]) outperformed the model-based method, achieving a speedup ratio of 2.86×2.86$\times$2.86 ×. Meanwhile, these results also underscores the potential benefits of integrating SAM-Decoding with different auxiliary strategies, especially in tasks where retrieval can effectively support the decoding process.
Table 2 presents the experimental results on the HumanEval, MBPP, and HAGRID datasets when using Vicuna-7B-v1.3. For the coding datasets HumanEval and MBPP, we constructed static suffix automata based on the python-code-instructions-18k dataset. For the contextual question answering dataset HAGRID, we built a static suffix automaton using its own training set.
In the case of coding tasks, SAM-Decoding significantly accelerates the model-free method. When SAM-Decoding is combined with Token Recycling, the speedup ratios improve from 2.78×2.78$\times$2.78 × and 2.83×2.83$\times$2.83 × to 2.94×2.94$\times$2.94 × and 2.87×2.87$\times$2.87 × on the HumanEval and MBPP datasets, respectively. Notably, SAM-Decoding demonstrates a more pronounced acceleration effect on HumanEval. This is attributed to the fact that HumanEval often features longer, more semantically rich variable names and includes additional annotations in the input prompts. These conditions are favorable for retrieval-based methods. Conversely, for model-based methods like EAGLE, integrating SAM-Decoding tends to decrease the average number of accepted tokens and the generation throughput.
Turning to the contextual question answering dataset HAGRID, this task is highly conducive to retrieval methods, and SAM-Decoding effectively enhances inference speed. When SAM-Decoding is paired with Token Recycling, EAGLE, and EAGLE2, their inference speeds on the HAGRID dataset increase from 1.93×1.93$\times$1.93 ×, 2.11×2.$1$, and 2.41×2.41$\times$2.41 × to $2.57$, $2.72$, and $2.81$, respectively. It is also evident that the model-free method SAM-Decoding[T] achieves a higher speedup on this dataset compared to the state-of-the-art model-based method EAGLE2.
In addition to Vicuna-7B, we also conducted experiments on more models, and the experimental results are shown in Appendix B.
Figure 2: The speedup ratio and mean accepted toknes of SAM-Decoding[T] under different $l_{\rm bias}$ and $l_{\rm threshold}$.
Figure 3: The speedup ratio and mean accepted toknes of SAM-Decoding[T] under different draft size.
Ablation Experiments. To further understand the contributions of various components of SAM-Decoding and the influence of different hyperparameters on inference speed, we conducted a series of ablation studies. Unless otherwise stated, these experiments were carried out on Spec-Bench using the vicuna-7B-v1.3 model, with SAM-Decoding[T] serving as the decoding strategy.
Firstly, we examined the effects of $l_{\rm bias}$ and $l_{\rm threshold}$ on inference speed through a grid search. These parameters control the preference for generating text from the generating text during retrieval and the preference for using retrieval results over the draft model when creating drafts. The findings are summarized in Figure 2. As shown, both the mean accepted tokens (MAT) and the speedup ratio increase with $l_{\rm bias}$ and $l_{\rm threshold}$ up to a point. Beyond a value of 5 for both parameters, these metrics start to decline.
Additionally, we investigated how the draft size utilized by SAM-Decoding affects inference speed. Figure 3 illustrates the mean accepted tokens and the inference speed of SAM-Decoding[T] at varying draft sizes. The data indicate that as the draft size grows, the mean number of accepted tokens and inference speed initially rise. However, once the draft size surpasses 40, these metrics begin to decline, and the decrease becomes more pronounced when the draft size reaches 70.
Table 3: The impact of different draft generation modules on inference speed.
Method Spec-Bench
#MAT Tokens/s Speedup
SAM-Decoding[T] 3.03 85.73 2.27×2.27$\times$2.27 ×
   w/o Static SAM 3.01 84.46 2.23×2.23$\times$2.23 ×
   w/o Dynamic SAM 2.68 76.14 2.01×2.01$\times$2.01 ×
   w/o Token Recycling 1.89 62.26 1.65×1.65$\times$1.65 ×
Finally, we investigated the impact of different modules within SAM-Decoding on inference speed. SAM-Decoding comprises three draft generation modules: the static suffix automaton, the dynamic suffix automaton, and the auxiliary speculative decoding module. We measured the inference speed of SAM-Decoding after removing each of these three modules individually. The results are presented in Table 3. From the experimental results, it is clear that each module contributes to the acceleration of the decoding process. Notably, the dynamic suffix automaton has a significantly greater impact compared to the static suffix automaton. This suggests that, in many cases, generating drafts from the dynamic context is more effective than retrieving drafts from a pre-existing text corpus. Additionally, since not all scenarios are well-suited for retrieval-based methods, the auxiliary speculative decoding module plays a crucial role in enhancing the overall performance of SAM-Decoding.
For more ablation experiment results, please refer to Appendix C.
5 Related Work
Speculative Decoding. Speculative decoding is an approach that can significantly speed up large language models (LLMs) without compromising the quality of their outputs. This is achieved by increasing the number of tokens produced in each inference step while simultaneously decreasing the total number of inference steps required. The majority of speculative decoding techniques rely on smaller neural networks to create initial drafts during the inference process. These techniques are referred to as model-based speculative decoding methods. Early implementations of model-based speculative decoding, such as those described in Speculative Decoding Leviathan et al. (2023), primarily focused on generating draft sequences using pre-existing, smaller-scale LLMs. Subsequently, advancements like Medusa Cai et al. (2024) and SpecInfer Miao et al. (2024) introduced tree-based speculative methods and initiated the development of dedicated draft models tailored for speculative decoding. For instance, Medusa developed a set of specialized decoding heads designed to serve as draft models. More recently, EAGLE Li et al. (2024b, a) highlighted the significance of incorporating the hidden state of the model’s output into the speculative decoding process. In line with this insight, EAGLE trained a draft model that integrates the hidden state information from the model’s output, leading to more efficient acceleration across a range of tasks.
In contrast to model-based methods, certain approaches focus on generating drafts by leveraging the model’s recent prediction outcomes, such as the output from the decoding head. Examples of such methods include Lookahead Decoding Fu et al. (2024) and Token Recycling Luo et al. (2024). Lookahead Decoding keeps track of the $n$-grams of the generated text, whereas Token Recycling maintains the top $k$ probable next tokens for each tokens.
Moreover, some strategies generate drafts directly through the retrieval of previously generated texts or text corpora, utilizing $n$-gram matching. Notable among these are PLD Saxena (2023) and REST He et al. (2024). Our proposed SAM-Decoding method also employs retrieval for draft generation. However, it stands out due to its superior speed and accuracy, and it can be seamlessly integrated with other speculative decoding techniques.
Additionally, beyond the aforementioned methods, research also conducted on speculative decoding that relies either on the model itself Kou et al. (2024) or on sub-models within the larger architecture Elhoushi et al. (2024).
Efficient LLM Architecture. There is also work to improve the model’s inference speed from the perspective of model structure. This part of the work includes model distillation, quantization and pruning. Model distillation Sreenivas et al. (2024); Muralidharan et al. (2024) distills the knowledge of a large model into a small model, thereby speeding up inference while maintaining the model’s performance. Quantization Frantar et al. (2022); Xiao et al. (2023); Lin et al. (2024); Liu et al. (2024); Ashkboos et al. (2024b) reduces the number of bits required to store parameters and reduces the data transmission time from HBM to on-chip memory during inference, thereby achieving effective inference acceleration. Pruning Frantar and Alistarh (2023); Ashkboos et al. (2024a); Men et al. (2024); Chen et al. (2024); Hu et al. (2024); Sun et al. (2024); Zhang et al. (2024) is used to remove unimportant parameters in the model. For structured pruning, it can be combined with model distillation to train efficient small models, while semi-structured pruning can reduce the model’s memory access and computing overhead and improve the inference speed by combining special hardware, i.e., sparse tensor core.
6 Conclusion
In this work, we propose SAM-Decoding, an speculative decoding method via suffix automatons constructed from both generated text and text corpus. SAM-Decoding can efficiently retrieve drafts from retrieval sources, thereby accelerating inference. SAM-Decoding is also designed to seamlessly integrate with existing methods. Consequently, in scenarios where retrieval is not feasible, SAM-Decodingcan adaptively switch to alternative methods for draft generation. Experimental results demonstrate that when combined with state-of-the-art techniques, SAM-Decodingcan significantly enhance performance in multi-turn conversation, summarization, and retrieval-augmented generation tasks.
Appendix A Suffix Automaton
A.1 Construction Process of Suffix Automaton
Algorithm 3 introduces the construction (Build-SAM) and expansion process (Expand) of Suffix Automaton, where INIT_SAM function will create a suffix automaton that only contains the root node. And for the root node, the link attribute value is $-1$, the next attribute value is empty, the length attribute value is 00, and the min_endpos attribute value is 00.
Algorithm 3 Construction Process of Suffix Automaton
  function Expand-State
     Input: suffix automaton $S$, link $l$, next $n$, length $len$, position $p$
[…]
     $s.\text{link}=l$
     $s.\text{next}=n$
     $s.\text{length}=len$
     $s.\text{min\_endpos}=p$
     return $s$
  end function
  function Expand
     Input: suffix automaton $S$, token $t$
[…]
     $c=\text{Expand-State}(S,-1,\{\},l,l)$
     $p=S.\text{last}$
[…]
        $p.\text{next[t]}=c$
        $p=p.\text{link}$
     end while
     if $p=\textbf{None}$ then
        $c.\text{link}=S.\text{root}$
     else
[…]
        if p.length + 1 = q.length then
           $c.\text{link}=q$
        else
           $cl=\text{Expand-State}(S,,-1,\{\},-1,-1)$
           $cl.\text{link}=q.\text{link}$
           $cl.\text{next}=q.\text{next}$
           $cl.\text{length}=p.\text{length}+1$
[…]
           while $p\neq\textbf{None}$ and $p.\text{next[t]}=q$ do
              $p.\text{next[t]}=cl$
              $p=p.\text{link}$
           end while
           $q.\text{link}=c.\text{link}=cl$
        end if
     end if
     $S.\text{last}=c$
  end function
  function Build-SAM
     Input: token sequence $s$
[…]
     for $t$ in $s$ do
        $\text{Expand}(S,t)$
     end for
     return $S$
  end function
A.2 Time Complexity of State Transfer
In this section, we introduce the time complexity of state transfer of suffix automaton. Consider a suffix automaton $S$ with initial state $s_{0}$, which corresponds to the root node of the automaton (representing the empty string). Suppose that state $s_{0}$ undergoes transitions through a sequence of $L$ tokens $x=(x_{1},x_{2},\ldots,x_{L})$:
$s_{i}=\text{Transfer}(S,x_{i},s_{i-1}),\quad i\in\{1,2,\ldots,L\}.$
We aim to demonstrate that the average time complexity of each state transition is $O(1)$, while the worst-case time complexity is $O(L)$.
First, let us define the matching length associated with state $s_{i}$ as $l_{i}$. Given that each state transition can increase the length of the match by at most 1, it follows that $0\leq l_{i}\leq i$. Next, we introduce the concept of energy $\phi$ for each state $s_{i}$, defined as $\phi(s_{i})=l_{i}$. Let $c_{i}$ represent the time cost of the transition of the $i$-th state. We then define the amortized cost $\hat{c_{i}}$ as:
$\hat{c_{i}}=c_{i}+\phi(s_{i})-\phi(s_{i-1}).$
We can now express the total amortized cost over all transitions as:
$\sum_{i=1}^{L}\hat{c_{i}}$ $=\sum_{i=1}^{L}(c_{i}+\phi(s_{i})-\phi(s_{i-1}))$
$=\sum_{i=1}^{L}c_{i}+\phi(s_{L})-\phi(s_{0}).$
Since $\phi(s_{i})\geq 0$ and $\phi(s_{0})=0$, it follows that:
$\sum_{i=1}^{L}\hat{c_{i}}\geq\sum_{i=1}^{L}c_{i}.$
Next, we analyze the upper bound of $\hat{c_{i}}$. Each state transition involves moving through the link edge zero or more times, followed by a move through the next edge. Transitioning through the link edge incurs a cost of 1 but decreases the potential by at least 1. Conversely, transitioning through the next edge also incurs a cost of 1 but increases the potential by 1. Consequently, the amortized cost $\hat{c_{i}}$ is bounded above by 2, leading to:
$\sum_{i=1}^{L}\hat{c_{i}}\leq 2L.$
Thus, the average time complexity of state transitions is:
$\frac{\sum_{i=1}^{L}c_{i}}{L}\leq\frac{2L}{L}=2,$
which is $O(1)$. In the worst case, a single operation may require up to $l_{i}$ transitions through the link edge, followed by one transition through the next edge, resulting in a worst-case time complexity of $O(L)$.
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
Llama3-8B-Inst PLD 1.30$\times$ 1.12$\times$ 1.41$\times$ 1.03$\times$ 1.30$\times$ 1.53$\times$ $1.39$ 44.26 1.28$\times$
Token Recycling 1.92$\times$ 1.88$\times$ 1.85$\times$ 1.75$\times$ 2.24$\times$ 1.82$\times$ 2.76 66.42 1.91$\times$
SAM-Decoding[T] 2.09$\times$ 1.93$\times$ 2.04$\times$ 1.82$\times$ 2.32$\times$ 2.12$\times$ 2.63 85.73 2.05$\times$
EAGLE2 2.08$\times$ 1.95$\times$ 1.85$\times$ 1.80$\times$ 2.31$\times$ 1.87$\times$ 3.90 68.69 1.98$\times$
SAM-Decoding[E2] 2.23$\times$ 1.92$\times$ 1.93$\times$ 1.76$\times$ 2.23$\times$ 2.06$\times$ 3.91 70.40 2.03$\times$
Table 4: Speedup of SAM-Decoding compared to the baselines on Spec-Bench.
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Llama3-8B-Inst PLD 1.30 42.39 1.18×1.18$\times$1.18 × 1.23 39.19 1.06×1.06$\times$1.06 × 1.50 45.15 1.34×1.34$\times$1.34 ×
Token Recycling 2.93 71.49 1.99×1.$9$ 2.89 74.05 2.02×2.02$\times$2.02 × 2.84 62.77 1.87×1.87$\times$1.87 ×
SAM-Decoding[T] 2.77 78.04 2.16×2.16$\times$2.16 × 2.65 76.76 2.09×2.09$\times$2.09 × 2.70 66.76 1.99×1.$9$
EAGLE2 4.74 85.58 2.37×2.37$\times$2.37 × 4.77 95.02 2.58×2.58$\times$2.58 × 3.97 63.30 1.86×1.86$\times$1.86 ×
SAM-Decoding[E2] 4.76 91.50 2.54×2.54$\times$2.54 × 4.34 89.15 2.43×2.43$\times$2.43 × 3.93 67.94 2.02×2.02$\times$2.02 ×
Table 5: Speedup of SAM-Decoding compared to the baselines on HumanEval, MBPP, and HAGRID.
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
Vicuna-13B PLD 1.61$\times$ 1.10$\times$ 2.36$\times$ 1.11$\times$ 1.69$\times$ 1.80$\times$ 1.66 33.89 1.59$\times$
Token Recycling 2.03$\times$ 1.84$\times$ 2.07$\times$ 1.83$\times$ 2.42$\times$ 1.84$\times$ 2.81 42.74 2.01$\times$
SAM-Decoding[T] 2.36$\times$ 1.80$\times$ 2.63$\times$ 1.83$\times$ 2.49$\times$ 2.22$\times$ 2.91 47.27 2.22$\times$
EAGLE2 3.10$\times$ 2.15$\times$ 2.58$\times$ 2.38$\times$ 3.19$\times$ 2.33$\times$ 4.42 56.06 2.63$\times$
SAM-Decoding[E2] 3.27$\times$ 2.12$\times$ 2.89$\times$ 2.34$\times$ 3.12$\times$ 2.54$\times$ 4.51 57.88 2.72$\times$
Table 6: Speedup of SAM-Decoding compared to the baselines on Spec-Bench.
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Vicuna-13B PLD 1.54 32.06 1.44×1.$4$ 1.40 29.43 1.30×1.30$\times$1.30 × 1.90 43.38 2.15×2.15$\times$2.15 ×
Token Recycling 2.79 46.03 2.07×2.07$\times$2.07 × 2.85 48.89 2.16×2.16$\times$2.16 × 2.90 40.97 2.03×2.03$\times$2.03 ×
SAM-Decoding[T] 2.79 50.87 2.28×2.28$\times$2.28 × 2.83 52.14 2.30×2.30$\times$2.30 × 2.99 48.33 2.40×2.40$\times$2.40 ×
EAGLE2 5.15 77.85 3.49×3.49$\times$3.49 × 5.26 82.13 3.63×3.63$\times$3.63 × 4.24 52.28 2.59×2.59$\times$2.59 ×
SAM-Decoding[E2] 4.97 73.78 3.31×3.31$\times$3.31 × 4.80 73.94 3.27×3.27$\times$3.27 × 4.41 56.17 2.78×2.78$\times$2.78 ×
Table 7: Speedup of SAM-Decoding compared to the baselines on HumanEval, MBPP, and HAGRID.
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
Vicuna-33B PLD 1.50$\times$ 1.07$\times$ 2.06$\times$ 1.09$\times$ 1.59$\times$ 1.51$\times$ 1.65 13.33 1.46$\times$
Token Recycling 2.10$\times$ 1.84$\times$ 2.19$\times$ 1.88$\times$ 2.42$\times$ 1.92$\times$ 2.70 18.8 2.06$\times$
SAM-Decoding[T] 2.31$\times$ 1.79$\times$ 2.53$\times$ 1.90$\times$ 2.48$\times$ 2.06$\times$ 2.68 19.87 2.18$\times$
EAGLE2 3.29$\times$ 2.31$\times$ 2.73$\times$ 2.51$\times$ 3.65$\times$ 2.46$\times$ 4.06 25.86 2.83$\times$
SAM-Decoding[E2] 3.36$\times$ 2.24$\times$ 2.88$\times$ 2.41$\times$ 3.43$\times$ 2.50$\times$ 4.08 25.60 2.81$\times$
Table 8: Speedup of SAM-Decoding compared to the baselines on Spec-Bench.
Model Method HumanEval MBPP HAGRID
#MAT Tokens/s Speedup #MAT Tokens/s Speedup #MAT Tokens/s Speedup
Vicuna-33B PLD 1.58 14.18 1.51×1.51$\times$1.51 × 1.34 11.82 1.30×1.30$\times$1.30 × - - -
Token Recycling 2.64 19.64 2.09×2.09$\times$2.09 × 2.76 20.92 2.19×2.19$\times$2.19 × - - -
SAM-Decoding[T] 2.73 22.44 2.39×2.39$\times$2.39 × 2.65 21.80 2.28×2.28$\times$2.28 × 2.60 19.74 2.26×2.26$\times$2.26 ×
EAGLE2 3.53 28.18 3.00×3.$0$ 5.04 37.35 3.91×3.91$\times$3.91 × 3.84 24.28 2.78×2.78$\times$2.78 ×
SAM-Decoding[E2] 3.61 29.49 3.14×3.14$\times$3.14 × 4.70 34.21 3.58×3.58$\times$3.58 × 3.79 24.45 2.80×2.80$\times$2.80 ×
Table 9: Speedup of SAM-Decoding compared to the baselines on HumanEval, MBPP, and HAGRID.
Model Method Multi-turn Conversation Translation Summarization Question Answering Mathematical Reasoning Retrieval-aug. Generation #Mean Accepted Tokens Tokens/s Overall
Vicuna-7B PLD 1.61$\times$ 1.56$\times$ 0.99$\times$ 2.47$\times$ 1.11$\times$ 1.51$\times$ 1.75 78.26 1.56$\times$
Token Recycling 2.08$\times$ 1.76$\times$ 1.97$\times$ 1.85$\times$ 2.35$\times$ 1.76$\times$ 2.82 98.39 1.96$\times$
SAM-Decoding[T] 2.62$\times$ 1.82$\times$ 2.92$\times$ 2.09$\times$ 2.60$\times$ 2.21$\times$ 3.02 119.21 2.38$\times$
EAGLE2 2.66$\times$ 1.76$\times$ 2.18$\times$ 2.03$\times$ 2.63$\times$ 1.97$\times$ 4.34 110.56 2.21$\times$
SAM-Decoding[E2] 3.19$\times$ 1.97$\times$ 2.86$\times$ 2.28$\times$ 2.84$\times$ 2.32$\times$ 4.52 129.36 2.58$\times$
Table 10: Speedup of SAM-Decoding on A100 GPU compared to the baselines on Spec-Bench.
Appendix B Additional Experiment Results
In this section, we present the results of the experiment on Llama3-8B-instruct.
Tables 4 and 5 present the speedup ratios of SAM-Decoding compared to baseline methods across the Spec-Bench, HumanEval, MBPP, and HAGRID datasets, utilizing the Llama3-8B-instruct model. In particular, SAM-Decoding , when paired with Token Recycling (SAM-Decoding[T]), brings speedups on all tasks and achieves the highest average speedup. Specifically, SAM-Decoding enhances the speedup ratio of Token Recycling from 1.92$\times$, 1.85$\times$, and 1.82$\times$ to 2.09$\times$, 2.04$\times$, and 2.12$\times$ for multi-turn conversation, summarization, and retrieval-augmented generation tasks, respectively. This improvement raises the overall speedup ratio of token recycling in the Spec-Bench dataset from 1.91$\times$ to 2.05$\times$. On the HumanEval and MBPP datasets, SAM-Decoding increases the speedup ratio of Token Recycling from 1.80$\times$ and 2.04$\times$ to 2.27$\times$ and 2.37$\times$, respectively. For the HAGRID dataset, the integration of SAM-Decoding with Token Recycling increases the speedup ratio from 1.87$\times$ to 1.99$\times$. Furthermore, beyond enhancing Token Recycling, SAM-Decoding also amplifies the performance gains of EAGLE2 in multi-turn conversation, summarization, retrieval-augmented generation, and context Q&A tasks. The speedup ratios were increased from 2.08$\times$, 1.85$\times$, 1.87$\times$, and 1.86$\times$ to 2.23$\times$, 1.93$\times$, 2.06$\times$, and 2.02$\times$ respectively.
Appendix C Additional Ablation Experiments
Figure 4: the percentage of inference time of different modules in SAM-Decoding.
Figure 5: the percentage of usage times and mean accept tokens of different draft modules.
In this section, we present additional ablation experiments, including the percentage of inference time of different modules in the decoding process of SAM-Decoding, and the percentage of drafts provided by different draft modules in SAM-Decoding.
The inference process of SAM-Decoding is divided into five stages: prefill, draft generation, decoding, verification, and updating. During the prefill stage, the model processes the input prompt to establish an initial state. In the draft generation stage, a draft is produced based on this initial state. The decoding stage involves the model further processing this draft. Next comes verification, where the correct parts of the draft are evaluated based on the information processed during the decoding stage. Finally, the update phase modifies the state of the model based on the valid parts of the draft. Figure 4 illustrates the proportion of time each stage consumes within the SAM-Decoding[T] process base on Spec-Bench. As shown, the decoding stage takes up the largest portion of time, accounting for 65.4% of the entire process. This is followed by the verification stage, which occupies 23.4% of the total time. The updating stage requires 6.3% of the time, whereas the draft generation stage contributes only 0.6% to the overall duration. Additionally, the prefill stage comprises 4.2% of the total processing time.
Figure 5 shows the usage times of different draft modules of SAM-Decoding[T] on Spec-Bench and the corresponding average draft accept length. It can be seen that in 85.96% of the cases, due to insufficient matching length, we generate drafts based on the auxiliary method, corresponding to an average accept length of 2.51, while in the remaining 11.59% and 2.45% of the cases, the dynamic suffix automaton and static suffix automaton are used to generate drafts, corresponding to average accept lengths of 6.57 and 3.39, respectively.